# VISCO: Benchmarking Fine-Grained Critique and Correction Towards Self-Improvement in Visual Reasoning

### v1

@@ 3.1.1 Critique
Detailed illustration of VISCore calculation will be in Appendix A.2.
@@ 3.2 Dataset Construction
The construction of VISCO dataset consists of four steps: task input collection, response collection, response filtering, and critique collection. More examples of each step and the final dataset will be in Appendix A.1.
@@
Critique collection. We train three expert human annotators to annotate critique $\mathcal{C}$ at three granularity. We first ask human annotators to verify the automatically generated answer critique $\mathcal{C}_{a}$. Then, we ask the annotators to judge the correctness of each step $s_{i}$, resulting in annotations of $\mathcal{C}_{s}$. This annotation achieves a substantial inter-annotator agreement of 64.76 Cohen’s Kappa. For disagreed data points, we employ a secondary annotation step where the annotators review each others’ annotations and engage in discussions to settle disagreements. Finally, we ask each annotator to write a natural language explanation for each incorrect step to collect $\mathcal{C}_{e}$, resulting in three references for each step. When evaluating the model-generated critique explanation, we compare it against each of the three annotations and take the average score, leading to a more accurate evaluation.
@@ 4.1 Experimental Settings
We evaluate 17 open-source general-purpose LVLMs from 7 families, including DeepSeek-VL [26], LLaVA [23], InternVL [7], Qwen-VL [39], Molmo [9], NVLM [8], and Llama-3.2. We further evaluate two open LVLMs trained to provide critique, Prometheus-Vision [15] and LLaVA-Critic [42]. For proprietary LVLMs, we evaluate three leading LVLMs, GPT-4o, Claude-3.5-Sonnet, and Gemini-Pro. As in Section 3.1, we use VISCore to evaluate critique and correction gain to evaluate correction. When measuring explanation-level critique F1, GPT-4o is adopted for LLM-assisted evaluation. For the critique task, we further establish a human baseline performed by trained expert annotators. For the correction task, we evaluate two settings, correction based on human-annotated ground truth critique and correction based on self-generated critique to simulate the self-improvement setting. More details about the experimental settings are in Appendix B.
@@ 5 LookBack as Improved Baseline
In this section, we present an initial step toward enhancing the critique capabilities of LVLMs. As analyzed in Section 4.2, the failure to critique visual perception and the reluctance to “say no” are two major bottlenecks for effective critique. To prevent LVLMs from blindly agreeing with the visual perception in the model response, our proposed LookBack framework forces LVLMs to explicitly check the image and verify the visual information before performing the critique.
As shown in Figure 17, for each step in the CoT, LookBack extracts a set of atomic information necessary for verification in the model response. LookBack then answers each question individually based on the image, and accordingly produces the final critique. This framework better simulates the human process of verifying and critiquing visual reasoning, as humans tend to revisit the image multiple times to iteratively validate every single detail against the image. We additionally include the original question $q$ into the set of questions to be answered. While the LVLMs might not fully succeed in answering the original question, it proves beneficial to propose a candidate solution before performing critique. Detailed algorithm will be shown in Alg. 1 in Appendix D.
@@ 6 Conclusions
This paper presents VISCO, a novel benchmark designed to evaluate LVLMs in terms of critique and correction, two fundamental capabilities for self-improvement. VISCO evaluates fine-grained critique that identifies the correctness of each reasoning step in a chain-of-thought and further provides natural language explanations. We curate the dataset from a diverse range of 18 datasets across 8 tasks and extensively evaluate 24 LVLMs. Our experiments show while LVLMs perform well in correction given high-quality critique, the effective generation of critique is the crucial bottleneck for self-improvement. We further investigate the model performance on critique tasks and identify three main issues contributing to LVLMs’ critique failures: failure to critique visual perception, reluctance to ‘say no”, and the exaggerated assumption of error propagation. To address these limitations, we designed the LookBack strategy, which can be combined with leading LVLMs to enhance the critique capability by up to 13.5 and further benefit the correction performance by up to 11.5. This work serves as an initial exploration into self-improvement of visual reasoning, aiming to inspire future research in this direction.
@@
Appendix A Additional Examples
A.1 Dataset Examples
@@
A.2 Examples for Calculating VISCore
Figure 25 illustrates the calculation of VISCore as the metric for critique task. Specifically, the calculation of explanation-level F1 requires matching model-generated critique explanations against human-annotated explanations using LLMs. We use GPT-4o for this evaluation, and the prompt is shown in Figure 26. To evaluate the reliability of this evaluation, we manually compare 100 model-generated explanations against human explanations, and evaluate the agreement between our manual evaluation and LLM evaluation. We observe a high agreement of 0.80 accuracy and 0.61 Cohen’s Kappa, validating the robustness of this evaluation approach.
A.3 Examples for Calculating Correction Gain
Figure 23 illustrates the calculation of correction gain as the metric for correction task.
Appendix B Experimental Details
We evaluate 26 open LVLMs and 3 proprietary LVLMs. For the open LVLMs, we run the inference with fast serving frameworks. Specifically, we evaluate Qwen2-VL, Molmo, Llama-3.2 and NVLM with vllm, evaluate LLaVA-OV and LLaVA-Critic with sglang, and evaluate InternVL2, DeepSeek-VL, LLaVA-v1.5, LLaVA-v1.6, Qwen-VL and Prometheus-Vision with lmdeploy. We set the sampling temperature as 0.7. The prompt for critique and correction are in Figure 27 and 28. We ask one of the trained annotators to establish the human baseline on a randomly selected subset of 265 data points. To reduce annotation costs, we provide the ground truth answers to the annotator. The 100% answer-level F1 shows that the ground truth answers are verified by the annotator to be correct.
Appendix C Additional Results
@@
Appendix D Details on LookBack

### v2

@@ 3.1.1 Critique
More details for VISCore design are in Appendix B.
@@ 3.2 Dataset Construction
The construction of VISCO dataset consists of four steps: task input collection, response collection, response filtering, and critique collection. More examples of each step and the final dataset are in Appendix A.
@@
Critique collection. We recruit three expert annotators (computer science students with relevant AI experience) to annotate critique $\mathcal{C}$. The first 30 examples are used for training the annotators and aligning annotation criteria; these are not included in the final dataset. During the official annotation process, annotators first verify the automatically generated answer critique $\mathcal{C}_{a}$, and then assess the correctness of each step $s_{i}$, resulting in annotations of $\mathcal{C}_{s}$. This annotation achieves a substantial inter-annotator agreement of 64.76 Cohen’s Kappa. For disagreed data points, we hold discussion sections every 50 samples where the annotators review each others’ annotations and settle disagreements. Finally, we ask each annotator to write a natural language explanation for each incorrect step to collect $\mathcal{C}_{e}$, resulting in three references for each step. When evaluating model-generated critique explanation, we compare it against each of the three annotations and take the average score, leading to more accurate evaluation.
@@ 4.1 Experimental Settings
We evaluate 18 open-source general-purpose LVLMs including LLaVA [23] and Llama-3.2, and three leading proprietary LVLMs including GPT-4o, Claude-3.5-Sonnet, and Gemini-Pro. We further evaluate two open LVLMs trained to provide critique, Prometheus-Vision [15] and LLaVA-Critic [42]. As in Section 3.1, we use VISCore to evaluate critique and correction gain to evaluate correction, adopting GPT-4o for LLM-assisted evaluation. For the critique task, we further establish a human baseline performed by trained expert annotators. For the correction task, we evaluate two settings, correction based on human-annotated ground truth critique and correction based on self-generated critique to simulate the self-improvement setting. More experimental details are in Appendix C.
@@ 5 LookBack as Improved Baseline
In this section, we present an initial step toward enhancing the critique capabilities of LVLMs. As in Section 4.2, the failure to critique visual perception and the reluctance to “say no” are two major bottlenecks for effective critique. To prevent LVLMs from blindly agreeing with the visual perception in the model response, we propose to explicitly check the image and verify the visual information before performing the critique.
An illustration of our proposed LookBack is shown in Figure 17, with the formal algorithm in Alg. 1. LookBack iteratively analyzes each step in the CoT, extracting atomic information and converting it into question format (L3). To ensure thoroughness, it incorporates the original question $q$ into the question set $\mathcal{Q}$ (L6). The extracted information is then verified by prompting these questions against the original image $I$ (L9), and the generated answers are synthesized into a holistic critique (L12). By breaking down complex reasoning into verifiable units and iteratively revisiting the image for validation, LookBack better mirrors the human critique process, where examiners cross-reference individual claims with visual evidence.
@@ 6 Conclusions
This paper presents VISCO, a novel benchmark designed to evaluate LVLMs in terms of critique and correction, two fundamental capabilities for self-improvement. VISCO evaluates fine-grained critique that identifies the correctness of each reasoning step in a chain-of-thought and further provides natural language explanations. We curate the dataset from a diverse range of 18 datasets across 8 tasks and extensively evaluate 24 LVLMs. Our experiments show while LVLMs perform well in correction given high-quality critique, the effective generation of critique is the crucial bottleneck for self-improvement. We further investigate the model performance on critique tasks and identify three main issues contributing to LVLMs’ critique failures: failure to critique visual perception, reluctance to “say no”, and the exaggerated assumption of error propagation. To address these limitations, we designed the LookBack strategy, which can be combined with leading LVLMs to enhance the critique capability by up to 13.5 and further benefit the correction performance by up to 11.5. This work serves as an initial exploration into self-improvement of visual reasoning, aiming to inspire future research in this direction.
Acknowledgment
This research is based upon work supported by CISCO, U.S. DARPA ECOLE Program No. #HR00112390060, and OFFICE OF NAVAL RESEARCH Award #N00014-23-1-2780. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ Acknowledgment
Appendix A Dataset Examples
@@
Appendix B Metric Design Details
VISCore design. To verify the robustness of VISCore metric design, we conducted human evaluation on 200 data where annotators compare pairs of model generated critique and judge which is better. We find that VISCore correlates the best with human judgment ($\tau=0.58$), outperforming alternatives like arithmetic mean of F1 ($\tau=0.42$) and exlanation BLEU ($\tau=0.41$). This is because VISCore prioritizes explanation correctness, better aligning with human focus, and LLM-based evaluation of explanation is more effective than BLEU.
LLM evaluation. The calculation of explanation-level F1 requires matching model-generated critique explanations against human-annotated explanations using LLMs. We use GPT-4o for this evaluation, and the prompt is shown in Figure 26. To evaluate the reliability of this evaluation, we manually compare 100 model-generated explanations against human explanations, and evaluate the agreement between our manual evaluation and LLM evaluation. We observe a high agreement of 0.80 accuracy and 0.61 Cohen’s Kappa, validating the robustness of LLM evaluation.
Metric calculation examples. To better illustrate the metric calculation, we present two examples in Figure 25 and 23, illustrating the calculation of VISCore for critique task and correction gain for correction task respectively.
Appendix C Experimental Details
We evaluate 27 open LVLMs and 3 proprietary LVLMs. For the open LVLMs, we evaluate models from 7 families: DeepSeek-VL [26], LLaVA [23], InternVL [7], Qwen-VL [39], Molmo [9], NVLM [8], Llama-3.2, and MiniCPM [45]. Specifically, we run the inference with fast serving frameworks. Specifically, we evaluate Qwen2-VL, Molmo, Llama-3.2, NVLM and MiniCPM with vllm, evaluate LLaVA-OV and LLaVA-Critic with sglang, and evaluate InternVL2, DeepSeek-VL, LLaVA-v1.5, LLaVA-v1.6, Qwen-VL and Prometheus-Vision with lmdeploy. We set the sampling temperature as 0.7. For LookBack, we use the same set of hyperparameters. The prompt for critique and correction are in Figure 27 and 28.
For the human baseline, we ask one of the trained annotators to establish the human baseline on a randomly selected subset of 265 data points. To reduce annotation costs, we provide the ground truth answers to the annotator. The 100% answer-level F1 shows that the ground truth answers are verified by the annotator to be correct.
Appendix D Additional Results
@@
To measure the importance of CoT to our framework, we perform ablation study by evaluating the critique and correction performance without CoT in the original response. As shown below, LVLMs can effectively perform critique and correction bringing positive correction gains without CoT. However, CoT brings better performance especially for strong models like GPT-4o.
Model Critique Correction
w. CoT w/o CoT w. CoT w/o CoT
InternVL2-8B 37.1 47.3 5.4 5.1
Claude-3.5-Sonnet 61.8 60.3 25.6 25.6
GPT-4o 63.0 59.0 28.8 26.4
Table 5: Additional experiments for ablations of CoT.
Appendix E Details on LookBack
@@
Error cases of LookBack. We further evaluate the error cases of LookBack and identify two typical scenarios: (1) Errors in reasoning critique. LookBack is less effective in critiquing reasoning, as these often go beyond verifying atomic information against the image. This is reflected in the smaller performance boosts in reasoning tasks in Table 4. (2) Referral expression error in atomic information extraction. In an example involving four images, the reasoning discusses a blue car in the first image, but the extracted question refers to the car as the blue car without specifying the first image, resulting in critique errors.
Appendix F Limitations
The main limitation of this work is its task setting, which constrains the CoT to at most five reasoning steps, potentially affecting its generalization to complex reasoning or multi-turn interactions. While this work provides a valuable initial investigation into LVLM critique and correction, this limitation highlights opportunities for future research.